# The occurrence of small, short-period planets younger than 200 Myr with TESS

### 05 Mar 2024 v1

@@ 7.2 Comparison to Hydrogen-helium dominated planet evolution models
Mapping the distribution of young planets has the potential to disentangle otherwise degenerate modes of planet formation. Some super-Earths and Neptune-sized planets may have formed in close-in volatile- and gas-depleted regions of the protoplanet disk (e.g. Lee et al., 2014), others may be composed of water-rich envelopes and were formed further out (e.g. Zeng et al., 2019; Luque & Pallé, 2022). These different formation modes produce planets of similar bulk physical properties in the mature planet distribution. In the water-rich scenario, the planets may have formed in the outer ice-rich protoplanet disk, migrating inward to their current locations (e.g. Mordasini et al., 2009). These planets may have silicate cores, icy mantles, and steam atmospheres (e.g. Rogers & Seager, 2010; Zeng et al., 2019).
@@
In the alternate scenario, the planets may be born with a thick hydrogen-helium envelope surrounding an iron-silicate core in the volatile-depleted inner regions of the disk (e.g. Lee et al., 2014). Some of these hydrogen-helium dominated planets may undergo runaway mass loss early in their evolution driven by boil-off (Owen & Wu, 2016; Ginzburg et al., 2016), photoevaporation (Lopez & Fortney, 2013; Owen & Wu, 2013) and core-powered (Ginzburg et al., 2018; Gupta & Schlichting, 2019) mechanisms. Owen & Schlichting (2023) suggested that a combination of these mechanisms is likely acting with timescales from Myrs to Gyrs, although the details of this timeline require further study.
@@
These different formation pathways differ most drastically early in their evolution. In particular, hydrogen-helium dominated atmospheres are expected to experience rapid contraction within the first few hundred million to billion years of evolution (e.g. Lopez et al., 2012; Chen & Rogers, 2016), while steam-dominated atmospheres experience very mild radius contraction as they cool (Lopez et al., 2012; Lopez, 2017), and our young planet population should resemble that of the Kepler distribution. Furthermore, atmospheric escape mechanisms should be evident with an increased frequency of planets around younger stars. As some of the planets have their atmospheres stripped and transition to below the radius valley, this frequency should decrease with time.
@@
For each star in the underlying SPOC and QLP stellar samples presented in Table 1, we forward-model 100 planets, assuming the underlying core mass, initial atmospheric mass fraction, core composition and orbital period distributions inferred in Rogers & Owen (2021). The synthesized planets have ages corresponding to that estimated for each respective association as listed in Table 1. We then randomly draw planets from this sample to match the Kepler occurrence rates, as done in Rogers & Owen (2021). This population, therefore represents a sample of young planets about our observed stellar sample, under the photoevaporation model, that would eventually evolve to reproduce the Kepler sample of Gyr old super-Earths and mini-Neptunes.
We compare the predicted yield to our QLP and SPOC planet yields in Figure 7. The models are in agreement at the level, showing the derived occurrence rates in this work are consistent with the contraction of hydrogen-dominated atmospheres undergoing photoevaporative escape (Rogers & Owen, 2021). Of note, one can see that the occurrence rate of super-Neptunes in the Rogers & Owen (2021) models is increased when compared to the Kepler sample in Figure 9. This is because the young mini-Neputunes in this model host significant hydrogen-helium atmospheres that are inflated to super-Neptune sizes for ages Myrs. While the results are generally inconsistent with non-contracting steam-based atmospheres (see Figure 7), we leave model comparisons of evolving hydrogen vs. high mean-molecular weight (e.g. steam) atmospheres for future work.
@@ 8 Conclusion
We identify a total of 23 planets, including 4 new planet candidates in the SPOC and QLP data previously not identified as TOIs. Initial ground-based follow-up rule out any obvious astrophysical false positives.
@@ Acknowledgements
We respectfully acknowledge the traditional custodians of the lands on which we conducted this research and throughout Australia. We recognize their continued cultural and spiritual connection to the land, waterways, cosmos and community. We pay our deepest respects to all Elders, present and emerging, and the people of the Giabal, Jarowair, and Kambuwal nations, upon whose lands this research was conducted.
SV thanks the support of the Fulbright U.S. Student Program, which is sponsored by the U.S. Department of State and Australian-American Fulbright Commission. The contents of this paper are solely the responsibility of the author and do not necessarily represent the official views of the Fulbright Program, the Government of the United States, or the Australian-American Fulbright Commission. GZ thanks the support of the Australian Research Council projects DE210101893 and FT230100517. CH thanks the support of the ARC DECRA project DE200101840.
@@
Appendix A Inverted Light Curve Test
False alarms are particularly prevalent at lower signal-to-noises. Kunimoto et al. (2023) illustrates the presence of an excess of false alarms from BLS searches of TESS light curves at periods related to both astrophysical and instrumental signatures. Specifically, Kunimoto et al. (2023) found build-ups of BLS false-alarm detections at aliases of a TESS orbit, days. As this work performs a BLS search for planets with periods on the order of a TESS orbit and 0.5 a TESS orbit, we perform a blind planet search on both inverted and non-inverted light curves to test the reliability of our detections.
We randomly sample 1000 stars in both our QLP and SPOC parent population to quantify our false alarm rate. We ensure that both subsamples include the light curves of all planet hosts detected in this work (see Table 2). For each individual light curve, we perform a random draw to determine if the light curve is to be inverted. For stars selected to have their light curves inverted, we follow Thompson et al. (2018), multiplying the zero-mean flux by negative one to invert the light curve. Each light curve is run through our planet detection pipeline (see Section 4).
From our test, no signals from our inverted light curve sample were wrongly classified as planet candidates. Of the positive light curves, all planet candidates were successfully recovered. In the QLP test sample, all the planet hosts were randomly selected to have their light curves inverted, none of which triggered a TCE. In the SPOC test sample, we identified TOI-2550.01 and DS Tuc Ab as planet candidates, neither of which had their light curves inverted. Additionally, HD 114082 b triggered a TCE, however, the period and radius are not within our search parameters and therefore was not identified as a planet candidate. The remaining planet hosts all had their light curves inverted. The inverted light curve of AU Mic triggered a TCE, but our vetting procedure rejected the event due to asymmetries in the detected possible transit events.
Appendix B Robustness of occurrence rate on sample assumptions
@@
B.1 Impact of RUWE threshold on occurrence rates
We investigate the impact of binarity on our derived occurrence rates by repeating our occurrence rate calculations without removing high RUWE stars. We collate our stellar parent population following Section 3. We find 8852 stars with TESS observations. Of which, 8505 were observed in the FFIs and 2243 were observed in the target pixel stamps.
@@ B.1 Impact of RUWE threshold on occurrence rates
B.2 Impact of false-positive rate assumptions on occurrence rates
To investigate the degree to which our adopted false-positive rate assumptions via TRICERATOPS impact our derived occurrence rates, we explore the bounding cases where (a) all planet candidates are counted as confirmed planets, and (b) only verified/published planets are included in our occurrence rate calculation and all planet candidates are treated as false-alarms. We calculate the occurrence rates following Section 6.
@@ Appendix C New Planetary Candidates
All new candidates have been submitted to the TESS Follow-Up Observing Program Working Group (TFOP WG) for ground-based follow-up observations for further confirmation and characterization. We also conducted photometric and spectroscopic ground-based follow-up observations for TIC 434398831 to allow us to rule out any false positive scenarios. Further, TIC 434398831 has been observed by the CHaracterising ExOPlanet Satellite (CHEOPS; Benz et al., 2021) as a part of AO-4 program PR230002 (PI Vach).
Precise planet radii and periods are key for calculating planetary occurrence rates. To properly characterize the newly identified planet candidates, we perform a global model to constrain the planetary parameters for the individual systems. We follow Zhou et al. (2021) and model the stellar properties and transit light curve parameters simultaneously. The free parameters for our global model are stellar mass, , stellar radius, , the orbital period, , the time of transit center, , the radius ratio, , and transit inclination , and eccentricity parameters and that influence the transit duration. The limb-darkening parameters are fixed to that interpolated from Claret (2017). The transits were calculated using the BATMAN (Kreidberg, 2015) implementation of the Mandel & Agol (2002) models. At each iteration, we interpolate the stellar parameters onto the MIST isochrone (Dotter, 2016) to estimate their magnitudes. These isochrone magnitudes are compared against that observed in the Gaia, (Gaia Collaboration et al., 2016), 2MASS, and WISE bands, corrected for the distance modulus via the Gaia parallax. We calculate a likelihood function incorporating the magnitudes and light curves. The best-fit parameters and posteriors are explored via a Markov chain Monte Carlo run via emcee (Foreman-Mackey et al., 2013). The derived planet radii and associated uncertainties are incorporated into our occurrence rate calculations (See Section 6). We present our best-fit values for the new planet candidates identified in this work below.
@@
TIC 150070085 was observed by TESS in Sectors 20 and 47, receiving both two-minute target pixel stamps and FFI observations. TIC 150070085.01 was identified in both our SPOC and QLP planet search with a signal-to-pink noise of 8.2 and 9.5 respectively. Our vetting diagnostic figure for TIC 150070085.01 is shown in Figure 10. We utilize the two-minute cadence observations for our global model.
@@
TIC 88785435 was observed by TESS in Sectors 11 and 38 in the FFIs. TIC 88785435.01 was identified in our QLP planet search with a signal-to-pink noise of 9.47. Figure 11 shows our vetting diagnostic figure for TIC 88785435.01.
@@
TIC 434398831 (Vach et al. in prep) received FFI observations over TESS sectors 6, 33, 43, 44, and 45. Our planet search identifies two planet candidates, TIC 434398831.01 (signal-to-pink noise of 11.8) and TIC 434398831.02 (signal-to-pink noise of 15.7). Our best fit parameters for TIC 434398831.01 are , days, BJD, and days, and for TIC 434398831.02 are days, BJD, and days.
@@
Three transits of TIC 434398831.01 and two transits of TIC 434398831.02 have been observed by ESA’s CHEOPS mission (AO-4 program PR230002). Paired with extensive LCO ground-based follow-up of TIC 434398831.01 and .02, we were able to clear the field of nearby eclipsing binaries (NEBs) and rule out any false-positive scenarios. Therefore, we classify TIC 434398831.01 and .02 as confirmed planets in our occurrence rate calculations and adopt false-positive rates of 0 for both (see Table 2).
@@ Appendix D Rotation Analysis Plots
Our parent stellar population is a collation of multiple catalogs (see section 3) from various clustering studies (Gagné et al., 2018; Kounkel & Covey, 2019; Ujjwal et al., 2020; Moranta et al., 2022). To independently measure the field star contamination rate of our collated stellar population, we make use of the TESS FFI light curves to measure the rotation periods and variability signatures of stars in our survey with K. As we do not attempt to measure rotation periods longer than the duration of a single TESS orbit, we sample periods between 0.5-12 days using a Lomb-Scargle periodogram. Figure 13 shows examples of vetting figures that were used to classify rotational modulation light curves. We visually examine each light curve for signatures of rotational variability, and report measured rotation periods where possible in Table 1.

### 11 Apr 2024 v2

@@ 7.2 Comparison to Hydrogen-Helium-dominated Planet Evolution Models
Mapping the distribution of young planets has the potential to disentangle otherwise degenerate modes of planet formation. Some super-Earths and Neptune-sized planets may have formed in close-in volatile- and gas-depleted regions of the protoplanet disk (e.g. Lee et al., 2014), while others may be composed of water-rich envelopes and were formed further out (e.g. Zeng et al., 2019; Luque & Pallé, 2022). These different formation modes produce planets of similar bulk physical properties in the mature planet distribution. In the water-rich scenario, the planets may have formed in the outer ice-rich protoplanet disk, migrating inward to their current locations (e.g. Mordasini et al., 2009). These planets may have silicate cores, icy mantles, and steam atmospheres (e.g. Rogers & Seager, 2010; Zeng et al., 2019).
@@
In the alternate scenario, the planets may be born with a thick hydrogen-helium envelope surrounding an iron-silicate core in the volatile-depleted inner regions of the disk (e.g. Lee et al., 2014). Some of these hydrogen-helium-dominated planets may undergo runaway mass loss early in their evolution driven by boil-off (Ginzburg et al., 2016; Owen & Wu, 2016), photoevaporation (Lopez & Fortney, 2013; Owen & Wu, 2013) and core-powered (Ginzburg et al., 2018; Gupta & Schlichting, 2019) mechanisms. Owen & Schlichting (2023) suggested that a combination of these mechanisms is likely acting with timescales from Myrs to gigayears, although the details of this timeline require further study.
@@
These different formation pathways differ most drastically early in their evolution. In particular, hydrogen-helium-dominated atmospheres are expected to experience rapid contraction within the first few hundred million to billion years of evolution (e.g. Lopez et al., 2012; Chen & Rogers, 2016), while steam-dominated atmospheres experience very mild radius contraction as they cool (Lopez et al., 2012; Lopez, 2017), and our young planet population should resemble that of the Kepler distribution. Furthermore, atmospheric escape mechanisms should be evident with an increased frequency of planets around younger stars. As some of the planets have their atmospheres stripped and transition to below the radius valley, this frequency should decrease with time.
@@
For each star in the underlying SPOC and QLP stellar samples presented in Table 1, we forward model 100 planets, assuming the underlying core mass, initial atmospheric mass fraction, core composition, and orbital period distributions inferred in Rogers & Owen (2021). The synthesized planets have ages corresponding to that estimated for each respective association as listed in Table 1. We then randomly draw planets from this sample to match the Kepler occurrence rates, as done in Rogers & Owen (2021). This population, therefore, represents a sample of young planets about our observed stellar sample, under the photoevaporation model, which would eventually evolve to reproduce the Kepler sample of Gyr old super-Earths and mini-Neptunes.
We compare the predicted yield to our QLP and SPOC planet yields in Figure 7. The models are in agreement at the level, showing the derived occurrence rates in this work are consistent with the contraction of hydrogen-dominated atmospheres undergoing photoevaporative escape (Rogers & Owen, 2021). Of note, one can see that the occurrence rate of super-Neptunes in the Rogers & Owen (2021) models is increased when compared to the Kepler sample in Figure 9. This is because the young mini-Neputunes in this model host significant hydrogen-helium atmospheres that are inflated to super-Neptune sizes for ages Myrs. While the results are generally inconsistent with noncontracting steam-based atmospheres (see Figure 7), we leave model comparisons of evolving hydrogen vs. high mean-molecular weight (e.g. steam) atmospheres for future work.
@@ 8 Conclusion
We identify a total of 23 planets, including four new planet candidates in the SPOC and QLP data previously not identified as TOIs. Initial ground-based follow up rule out any obvious astrophysical false positives.
@@ Acknowledgements
We respectfully acknowledge the traditional custodians of the lands on which we conducted this research and throughout Australia. We recognize their continued cultural and spiritual connection to the land, waterways, cosmos, and community. We pay our deepest respects to all Elders, present and emerging, and the people of the Giabal, Jarowair, and Kambuwal nations, upon whose lands this research was conducted.
S.V. thanks the support of the Fulbright U.S. Student Program, which is sponsored by the U.S. Department of State and Australian-American Fulbright Commission. The contents of this paper are solely the responsibility of the author and do not necessarily represent the official views of the Fulbright Program, the Government of the United States, or the Australian-American Fulbright Commission. G.Z. thanks the support of the Australian Research Council projects DE210101893 and FT230100517. C.H. thanks the support of the ARC DECRA project DE200101840.
@@
Appendix A Inverted Light-curve Test
False alarms are particularly prevalent at lower S/Ns. Kunimoto et al. (2023) illustrate the presence of an excess of false alarms from BLS searches of TESS light curves at periods related to both astrophysical and instrumental signatures. Specifically, Kunimoto et al. (2023) found buildups of BLS false-alarm detections at aliases of the TESS orbit, days. As this work performs a BLS search for planets with periods on the order of a TESS orbit and 0.5 a TESS orbit, we perform a blind planet search on both inverted and noninverted light curves to test the reliability of our detections.
We randomly sample 1000 stars in both our QLP and SPOC parent population to quantify our false-alarm rate. We ensure that both subsamples include the light curves of all planet hosts detected in this work (see Table 2). For each individual light curve, we perform a random draw to determine if the light curve is to be inverted. For stars selected to have their light curves inverted, we follow Thompson et al. (2018), multiplying the zero-mean flux by to invert the light curve. Each light curve is run through our planet detection pipeline (see Section 4).
From our test, no signals from our inverted light-curve sample were wrongly classified as planet candidates. Of the positive light curves, all planet candidates were successfully recovered. In the QLP test sample, all the planet hosts were randomly selected to have their light curves inverted, none of which triggered a TCE. In the SPOC test sample, we identified TOI-2550.01 and DS Tuc Ab as planet candidates, neither of which had their light curves inverted. Additionally, HD 114082 b triggered a TCE, however, the period and radius are not within our search parameters and therefore was not identified as a planet candidate. The remaining planet hosts all had their light curves inverted. The inverted light curve of AU Mic triggered a TCE, but our vetting procedure rejected the event due to asymmetries in the detected possible transit events.
Appendix B Robustness of the Occurrence Rate on Sample Assumptions
@@
B.1 Impact of the Renormalized Unit Weight Error Threshold on the Occurrence Rates
We investigate the impact of binarity on our derived occurrence rates by repeating our occurrence rate calculations without removing high-RUWE stars. We collate our stellar parent population following Section 3. We find 8852 stars with TESS observations. Of which, 8505 were observed in the FFIs and 2243 were observed in the target pixel stamps.
@@ B.1 Impact of the Renormalized Unit Weight Error Threshold on the Occurrence Rates
B.2 Impact of the False-positive Rate Assumptions on the Occurrence Rates
To investigate the degree to which our adopted false-positive rate assumptions via TRICERATOPS impact our derived occurrence rates, we explore the bounding cases where (a) all planet candidates are counted as confirmed planets and (b) only verified/published planets are included in our occurrence rate calculation and all planet candidates are treated as false-alarms. We calculate the occurrence rates following Section 6.
@@ Appendix C New Planetary Candidates
All new candidates have been submitted to the TFOP WG for ground-based follow-up observations for further confirmation and characterization. We also conducted photometric and spectroscopic ground-based follow-up observations for TIC 434398831 to allow us to rule out any false-positive scenarios. Further, TIC 434398831 has been observed by the CHaracterising ExOPlanet Satellite (CHEOPS; Benz et al., 2021) as a part of AO-4 program PR230002 (PI: S. Vach).
Precise planet radii and periods are key for calculating planetary occurrence rates. To properly characterize the newly identified planet candidates, we perform a global model to constrain the planetary parameters for the individual systems. We follow Zhou et al. (2021) and model the stellar properties and transit light-curve parameters simultaneously. The free parameters for our global model are stellar mass, , stellar radius, , the orbital period, , the time of transit center, , the radius ratio, , and transit inclination , and eccentricity parameters and that influence the transit duration. The limb-darkening parameters are fixed to those interpolated from Claret (2017). The transits were calculated using the BATMAN (Kreidberg, 2015) implementation of the Mandel & Agol (2002) models. At each iteration, we interpolate the stellar parameters onto the MIST isochrone (Dotter, 2016) to estimate their magnitudes. These isochrone magnitudes are compared against that observed in the Gaia, (Gaia Collaboration et al., 2016), the Two Micron All Sky Survey, and Wide-field Infrared Survey Explorer bands, corrected for the distance modulus via the Gaia parallax. We calculate a likelihood function incorporating the magnitudes and light curves. The best-fit parameters and posteriors are explored via a Markov Chain Monte Carlo run via emcee (Foreman-Mackey et al., 2013). The derived planet radii and associated uncertainties are incorporated into our occurrence rate calculations (See Section 6). We present our best-fit values for the new planet candidates identified in this work below.
@@
TIC 150070085 was observed by TESS in Sectors 20 and 47, receiving both 2 minute target pixel stamps and FFI observations. TIC 150070085.01 was identified in both our SPOC and QLP planet search with a signal-to-pink noise ratio of 8.2 and 9.5 respectively. Our vetting diagnostic figure for TIC 150070085.01 is shown in Figure 10. We utilize the 2 minute cadence observations for our global model.
@@
TIC 88785435 was observed by TESS in Sectors 11 and 38 in the FFIs. TIC 88785435.01 was identified in our QLP planet search with a signal-to-pink noise ratio of 9.47. Figure 11 shows our vetting diagnostic figure for TIC 88785435.01.
@@
TIC 434398831 (S. Vach et al. 2024, in preparation) received FFI observations over TESS Sectors 6, 33, 43, 44, and 45. Our planet search identifies two planet candidates, TIC 434398831.01 (signal-to-pink noise ratio of 11.8) and TIC 434398831.02 (signal-to-pink noise ratio of 15.7). Our best-fit parameters for TIC 434398831.01 are , days, BJD, and days, and for TIC 434398831.02 they are days, BJD, and days.
@@
Three transits of TIC 434398831.01 and two transits of TIC 434398831.02 have been observed by ESA’s CHEOPS mission (AO-4 program PR230002). Paired with extensive LCO ground-based follow up of TIC 434398831.01 and .02, we were able to clear the field of NEBs and rule out any false-positive scenarios. Therefore, we classify TIC 434398831.01 and .02 as confirmed planets in our occurrence rate calculations and adopt false-positive rates of zero for both (see Table 2).
@@ Appendix D Rotation Analysis Plots
Our parent stellar population is a collation of multiple catalogs (see section 3) from various clustering studies (Gagné et al., 2018; Kounkel & Covey, 2019; Ujjwal et al., 2020; Moranta et al., 2022). To independently measure the field star contamination rate of our collated stellar population, we make use of the TESS FFI light curves to measure the rotation periods and variability signatures of stars in our survey with K. As we do not attempt to measure rotation periods longer than the duration of a single TESS orbit, we sample periods between 0.5 and 12 days using a Lomb-Scargle periodogram. Figure 13 shows examples of vetting figures that were used to classify rotational modulation light curves. We visually examine each light curve for signatures of rotational variability, and report the measured rotation periods where possible in Table 1.